How many liters of a 40% acid mixture and how many liters of a 10% acid mixture should be mixed to obtain 20 L of a 22% acid mixture?

Answer by Alan3354(69443)

Do it like this one. Only the numbers are different.
-----------
Amanda wants to make 8 gal. of a 20% saline solution by mixing together a 56% saline solution and a 8% saline solution. How much of each solution must she use?
================
e = amount of 8%
f = amount of 56%
---
e+ f = 8 (total solution)
8e + 56f = 20*8 (total saline)
---
e+ f = 8
e + 7f = 20
------------ Subtract
-6f = -12
f = 2
e = 6